OpenStudy (anonymous):

to find arc length is it the degree/360*2*pi*r?

OpenStudy (anonymous):

arclength s = r * theta

OpenStudy (anonymous):

To convert from degrees to radians:\[\theta \times \frac{\pi}{180}.\] From radians to degrees, you instead multiply by 180/pi. \[\theta = \frac{s}{r}\] where theta is the angle in radians, s is the arc length, and r is the radius. Solve for whatever variable you need, and out comes your value.